# A real-time digital twin of azimuthal thermoacoustic instabilities

### v1

@@ 3.3 Stochastic ensemble framework
where $\otimes$ is the dyadic product. (Because the forecast operator $\mathcal{F}$ is nonlinear, the Gaussian prior may not remain Gaussian after the model forecast, and $\mathbb{E}(\mathcal{F}({\bm{\mathit{\psi}}}))\neq\mathcal{F}(\bar{{\bm{\mathit{\psi}}}})$. However, the time between analyses $\Delta t_{\mathrm{d}}$ is assumed small enough such that the Gaussian distribution is not significantly distorted (Evensen, 2009; Yu et al., 2019b).)
Finally, we approximate the model bias (10) and measurement bias (18) with the ensemble statistics
${\bm{\mathit{b}}}$ $\approx{\bm{\mathit{d}}}^{\dagger}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}},$ (25)
${\bm{\mathit{b}}}_{d}$ $\approx{\bm{\mathit{d}}}^{\dagger}-\bar{{\bm{\mathit{d}}}}.$ (26)
Hence, the innovation of the ensemble mean can be approximated as
@@ 3.4 The bias-regularized ensemble Kalman filter
The objective function in a bias-regularized data assimilation framework contains three norms (Nóvoa et al., 2024)
$\mathcal{J}({\bm{\mathit{\psi}}}_{j})=$ $\left\|{\bm{\mathit{\psi}}}_{j}-{\bm{\mathit{\psi}}}_{j}^{\mathrm{f}}\right\|^{2}_{\bm{\mathsf{C}}^{\mathrm{f}^{-1}}_{\psi\psi}}+\left\|{{\bm{\mathit{y}}}}_{j}-{\bm{\mathit{d}}}_{j}\right\|^{2}_{\bm{\mathsf{C}}^{-1}_{dd}}+\gamma\left\|{\bm{\mathit{b}}}_{j}\right\|^{2}_{\bm{\mathsf{C}}^{-1}_{bb}},\quad\mathrm{for}\quad j=0,\dots,m-1$ (28)
where the superscript ‘f’ indicates ‘forecast’; the operator $\left\|\cdot\right\|^{2}_{\bm{\mathsf{C}}^{-1}}$ is the $L_{2}$-norm weighted by the semi-positive definite matrix ${\bm{\mathsf{C}}^{-1}}$; $\gamma\geq 0$ is a user-defined bias regularization factor; and ${\bm{\mathit{b}}}_{j}$ is the model bias of each ensemble member. For simplicity, we define the bias in the ensemble mean, such that ${\bm{\mathit{b}}}_{j}={\bm{\mathit{b}}}$ for all $j$ (which we estimate with an echo state network §4(Nóvoa et al., 2024). From left to right, the norms on the right-hand-side of (28) measure (1) the spread of the ensemble prediction, (2) the distance between the bias-corrected estimate and the observables, and (3) the model bias. The analytical solution of the bias-regularized ensemble Kalman filter (r-EnKF) in (29), which globally minimizes the cost function (28) with respect to ${\bm{\mathit{\psi}}}_{j}$, is
${\bm{\mathit{\psi}}}_{j}^{\mathrm{a}}$ $={\bm{\mathit{\psi}}}_{j}^{\mathrm{f}}+\bm{\mathsf{K}}_{\mathrm{r}}\left[\left(\mathbb{I}+\bm{\mathsf{J}}\right)\left({\bm{\mathit{d}}}_{j}+{\bm{\mathit{b}}}_{d}^{\mathrm{f}}-{\bm{\mathit{y}}}_{j}^{\mathrm{f}}\right)-\gamma\bm{\mathsf{C}}_{dd}\bm{\mathsf{C}}^{-1}_{bb}\bm{\mathsf{J}}{\bm{\mathit{b}}}^{\mathrm{f}}\right],\quad j=0,\dots,m-1$ (29a)
@@
$\bm{\mathsf{K}}_{\mathrm{r}}=\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}\left[\bm{\mathsf{C}}_{dd}+(\mathbb{I}+\bm{\mathsf{J}})\bm{\mathsf{M}}\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}(\mathbb{I}+\bm{\mathsf{J}})^{\mathrm{T}}+\gamma\bm{\mathsf{C}}_{dd}\bm{\mathsf{C}}^{-1}_{bb}\bm{\mathsf{J}}\bm{\mathsf{M}}\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}{\bm{\mathsf{J}}}^{\mathrm{T}}\right]^{-1},$ (29b)
@@
where ‘a’ stands for ‘analysis’, i.e, the optimal state of the assimilation; we assimilate each ensemble member with a different ${\bm{\mathit{d}}}_{j}\sim\mathcal{N}({\bm{\mathit{d}}},\bm{\mathsf{C}}_{dd})$ to avoid covariance underestimation in ensemble filters (Burgers et al., 1998); $\bm{\mathsf{K}}_{\mathrm{r}}$ is the regularized Kalman gain matrix; and $\bm{\mathsf{J}}=\mathrm{d}\,{\bm{\mathit{b}}}/\mathrm{d}\,\bm{\mathsf{M}}{\bm{\mathit{\psi}}}$ is the Jacobian of the bias estimator. Formulae 29 generalize the analytical solutions of Nóvoa et al. (2024) to the case of biased measurement. We prescribe $\bm{\mathsf{C}}_{dd}=\bm{\mathsf{C}}_{bb}$ because the model bias is defined in the observable space. We use $\gamma$ to tune the norm of the bias in (28) (Nóvoa et al., 2024). The optimal state and parameters are
@@
$\begin{bmatrix}{\bm{\mathit{\phi}}}_{j}^{\mathrm{a}}\\[10.00002pt] {\bm{\mathit{\alpha}}}_{j}^{\mathrm{a}}\end{bmatrix}=\begin{bmatrix}{\bm{\mathit{\phi}}}_{j}^{\mathrm{f}}\\[10.00002pt] {\bm{\mathit{\alpha}}}_{j}^{\mathrm{f}}\end{bmatrix}+\overbrace{\begin{bmatrix}\bm{\mathsf{C}}_{\phi q}^{\mathrm{f}}\\[10.00002pt] \bm{\mathsf{C}}_{\alpha q}^{\mathrm{f}}\end{bmatrix}\left\{\bm{\mathsf{C}}_{dd}+(\mathbb{I}+\bm{\mathsf{J}})\bm{\mathsf{C}}_{qq}^{\mathrm{f}}(\mathbb{I}+\bm{\mathsf{J}})^{\mathrm{T}}+\gamma\bm{\mathsf{J}}\bm{\mathsf{C}}_{qq}^{\mathrm{f}}{\bm{\mathsf{J}}}^{\mathrm{T}}\right\}^{-1}}^{\mathrm{Regularized\,Kalman\,gain,\,\bm{\mathsf{K}}_{\mathrm{r}}}}\dots$
@@
$\dots\Big{[}\left(\mathbb{I}+\bm{\mathsf{J}}\right)\overbrace{\left({\bm{\mathit{d}}}_{j}+{\bm{\mathit{b}}}_{d}^{\mathrm{f}}-{\bm{\mathit{y}}}_{j}^{\mathrm{f}}\right)}^{\mathrm{Corrected\,innovation}}-\gamma\bm{\mathsf{J}}{\bm{\mathit{b}}}^{\mathrm{f}}\Big{]}$ (30)
@@
The r-EnKF defines a ‘good’ analysis from a biased model if the unbiased state ${\bm{\mathit{y}}}$ matches the truth, and the model bias ${\bm{\mathit{b}}}$ is small relative to the truth. The underlying assumptions of this work are that (i) our low-order model is qualitatively accurate such that the model bias ${\bm{\mathit{b}}}^{\mathrm{f}}$ has a small norm; and (ii) the sensors are properly calibrated. In the limiting case when the assimilation framework is unbiased (i.e., ${\bm{\mathit{b}}}={\bm{\mathit{0}}}$, and ${\bm{\mathit{b}}}_{d}={\bm{\mathit{0}}}$), the r-EnKF (3.4) becomes the bias-unregularized EnKF (A).
4 Reservoir computing for inferring model and measurement biases
@@
To apply the r-EnKF (3.4) we must provide an estimate of the model bias and the measurement bias. We employ an echo state network (ESN) for this task. ESNs are suitable for real-time data assimilation because (i) they are recurrent neural networks, i.e., they are designed to learn temporal dynamics in data; (ii) they are based on reservoir computing, hence they are universal approximators (Grigoryeva and Ortega, 2018); (iii) they are general nonlinear auto-regressive models (Aggarwal, 2018); and (iv) training an ESN consists of solving a linear regression problem, which provides a global minimum without back propagation. In this work, we generalize the implementation of Nóvoa et al. (2024) to account for both the model and measurement biases (§4.1). We propose a training dataset generation based on time-series correlation (§4.2).
@@ 4.1 Echo state network architecture and state-space formulation
We propose an ESN that simultaneously estimates ${\bm{\mathit{b}}}^{\mathrm{f}}$ and ${\bm{\mathit{b}}}_{d}^{\mathrm{f}}$ (where ‘f’ indicates that the biases are forecast quantities, which are equivalent to the actual biases at convergence). Importantly, the biases ${\bm{\mathit{b}}}$ and ${\bm{\mathit{b}}}_{d}$ are not directly measurable in real-time, but we have access to the innovation, which is linked to the biases through (27). Once we have information on the innovation and the model bias (the measurement bias), we can estimate the measurement bias (the model bias) through Eq. 27.
@@
Figure 3a represents pictorially the proposed ESN at time $t_{k}$, with its three main components: (i) the input data, which are the mean analysis innovations, i.e, $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$; (ii) the reservoir, which is a high-dimensional state is characterized by the sparse reservoir matrix $\bm{\mathsf{W}}\in\mathbb{R}^{N_{r}\times N_{r}}$ and vector ${\bm{\mathit{r}}}_{k}\in\mathbb{R}^{N_{r}}$ ($N_{r}\gg N_{q}$ is the number of neurons in the reservoir states); and (iii) the outputs, which are the mean innovation and the model bias at the next time step, i.e., $\bar{{\bm{\mathit{i}}}}^{\mathrm{f}}_{k+1}$ and ${\bm{\mathit{b}}}^{\mathrm{f}}_{k+1}$. The inputs to the network are a subset of the output (i.e., the innovation only). The sparse input matrix $\bm{\mathsf{W}}_{\mathrm{in}}\in\mathbb{R}^{N_{r}\times(N_{q}+1)}$ maps the physical state into the reservoir, and the output matrix $\bm{\mathsf{W}}_{\mathrm{out}}\in\mathbb{R}^{N_{q}\times(N_{r}+1)}$ maps the reservoir state back to the physical state.
@@
Figure 3: Schematic representation of the proposed echo state network architecture for model bias and innovation prediction. (a) Compact architecture showing the open-loop and closed-loop configurations; and (b) unfolded architecture starting at an analysis step is at $t_{k}$, in which there is one open-loop step followed by a closed-loop forecast. In open loop the input to the ESN is the analysis mean innovation $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$, i.e., the difference between the raw acoustic pressure data and the analysis model estimate. The outputs from the ESN are (i) the innovation at the next time step $\bar{{\bm{\mathit{i}}}}^{\mathrm{f}}_{k+1}$ (which becomes the input to the next time step in closed loop configuration); and (ii) the model bias ${\bm{\mathit{b}}}^{\mathrm{f}}_{k+1}$, i.e., an estimate of the difference between the presumed acoustic state ${\bm{\mathit{d}}}_{k+1}^{\dagger}$ and the model estimate ${\bm{\mathit{q}}}_{k+1}$.
@@
Further, Figure 3a shows the two forecast settings of the echo state network: open-loop, which is performed when data on the innovations are available, and the closed-loop, in which the ESN runs autonomously using the output as the input in the next time step. Figure 3b shows the unfolded architecture starting at time $t_{k}$ when observations of the innovations are available (i.e., at the analysis step). At this time, the reservoir is re-initialized with the analysis innovations such that the input to the ESN is $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$. From this, the ESN estimates at the next time step $t_{k+1}$ the model bias and the innovation $\bar{{\bm{\mathit{i}}}}^{\mathrm{f}}_{k+1}$, which is used as initial condition in the subsequent forecast step. Mathematically, the equations that forecast the model bias and mean innovation in time are
$[{\bm{\mathit{b}}}_{k+1}^{\mathrm{f}};\bar{{\bm{\mathit{i}}}}^{\mathrm{f}}_{k+1}]$ $=\bm{\mathsf{W}}_{\mathrm{out}}\left[{\bm{\mathit{r}}}_{k+1};1\right]$
@@
${\bm{\mathit{r}}}_{k+1}$ $=\textrm{tanh}\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}\left[\bar{{\bm{\mathit{i}}}}^{\star}_{k}\odot{\bm{\mathit{g}}};\delta_{\mathrm{r}}\right]+\rho\bm{\mathsf{W}}{\bm{\mathit{r}}}_{k}\right),$ (31)
@@
where $\bar{{\bm{\mathit{i}}}}^{\star}=\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}$ in open loop and $\bar{{\bm{\mathit{i}}}}^{\star}=\bar{{\bm{\mathit{i}}}}^{\mathrm{f}}$ in closed loop; the ${\tanh(\cdot)}$ operation is performed element-wise; the operator $\odot$ is the Hadamard product, i.e., component-wise multiplication; and ${\bm{\mathit{g}}}=[g_{1};\dots;g_{2N_{q}}]$ is the input normalizing term with $g_{q}^{-1}=\max{({\bm{\mathit{u}}}_{q})}-\min{({\bm{\mathit{u}}}_{q})}$, i.e., $g_{q}^{-1}$ is the range of the $q^{\mathrm{th}}$ component of the training data $\bm{\mathsf{U}}$; $\delta_{\mathrm{r}}$ is a constant used for breaking the symmetry of the ESN (we set $\delta_{\mathrm{r}}=0.1$(Huhn and Magri, 2020). We define $\bm{\mathsf{W}}_{\mathrm{in}}$ and $\bm{\mathsf{W}}$ as sparse and randomly generated, with a connectivity of 3 neurons (see Jaeger and Haas, 2004, for details). We compute the matrix $\bm{\mathsf{W}}_{\mathrm{out}}$ during the training.

### v2

@@ 3.3 Stochastic ensemble framework
where $\otimes$ is the dyadic product. Because the forecast operator $\mathcal{F}$ is nonlinear, the Gaussian prior may not remain Gaussian after the model forecast, and $\mathbb{E}(\mathcal{F}({\bm{\mathit{\psi}}}))\neq\mathcal{F}(\bar{{\bm{\mathit{\psi}}}})$. However, the time between analyses $\Delta t_{d}$ is assumed small enough such that the Gaussian distribution is not significantly distorted (Evensen, 2009; Yu et al., 2019b).
Finally, we approximate the model bias as ${\bm{\mathit{b}}}\approx{\bm{\mathit{d}}}^{\mathrm{t}}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}$, thus, the sum of the biases is approximately equal to the mean of the innovations, i.e.,
@@ 3.4 The bias-regularized ensemble Kalman filter
The objective function in a bias-regularized ensemble data assimilation framework contains three norms (Nóvoa et al., 2024)
$\mathcal{J}({\bm{\mathit{\psi}}}_{j})=$ $\left\|{\bm{\mathit{\psi}}}_{j}-{\bm{\mathit{\psi}}}_{j}^{\mathrm{f}}\right\|^{2}_{\bm{\mathsf{C}}^{\mathrm{f}^{-1}}_{\psi\psi}}+\left\|{{\bm{\mathit{y}}}}_{j}-{\bm{\mathit{d}}}_{j}\right\|^{2}_{\bm{\mathsf{C}}^{-1}_{\mathrm{dd}}}+\gamma\left\|{\bm{\mathit{b}}}_{j}\right\|^{2}_{\bm{\mathsf{C}}^{-1}_{\mathrm{bb}}},\quad\mathrm{for}\quad j=0,\dots,m-1$ (24)
where the superscript ‘f’ indicates ‘forecast’; the operator $\left\|\cdot\right\|^{2}_{\bm{\mathsf{C}}^{-1}}$ is the $L_{2}$-norm weighted by the semi-positive definite matrix ${\bm{\mathsf{C}}^{-1}}$; $\gamma\geq 0$ is a user-defined bias regularization factor; and ${\bm{\mathit{b}}}_{j}$ is the model bias of each ensemble member. For simplicity, we define the bias in the ensemble mean, i.e., ${\bm{\mathit{b}}}_{j}={\bm{\mathit{b}}}$ for all $j$ (which we estimate with an echo state network in §4(Nóvoa et al., 2024). From left to right, the norms on the right-hand-side of (24) measure (1) the spread of the ensemble prediction, (2) the distance between the bias-corrected estimate and the observables, and (3) the model bias. The analytical solution of the bias-regularized ensemble Kalman filter (r-EnKF) in (25), which globally minimizes the cost function (24) with respect to ${\bm{\mathit{\psi}}}_{j}$, is
${\bm{\mathit{\psi}}}_{j}^{\mathrm{a}}$ $={\bm{\mathit{\psi}}}_{j}^{\mathrm{f}}+\bm{\mathsf{K}}_{r}\left[\left(\mathbb{I}+\bm{\mathsf{J}}\right)^{\mathrm{T}}\left({\bm{\mathit{d}}}_{j}+{\bm{\mathit{b}}}_{d}-{\bm{\mathit{y}}}_{j}^{\mathrm{f}}\right)-\gamma\bm{\mathsf{C}}_{dd}\bm{\mathsf{C}}^{-1}_{\mathrm{bb}}\bm{\mathsf{J}}^{\mathrm{T}}{\bm{\mathit{b}}}\right],\quad j=0,\dots,m-1$ (25a)
@@
$\bm{\mathsf{K}}_{r}=\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}\left[\bm{\mathsf{C}}_{dd}+(\mathbb{I}+\bm{\mathsf{J}})^{\mathrm{T}}(\mathbb{I}+\bm{\mathsf{J}})\bm{\mathsf{M}}\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}+\gamma\bm{\mathsf{C}}_{dd}\bm{\mathsf{C}}^{-1}_{\mathrm{bb}}\bm{\mathsf{J}}^{\mathrm{T}}\bm{\mathsf{J}}\bm{\mathsf{M}}\bm{\mathsf{C}}_{\psi\psi}^{\mathrm{f}}\bm{\mathsf{M}}^{\mathrm{T}}\right]^{-1},$ (25b)
@@
where ‘a’ stands for ‘analysis’, i.e, the optimal state of the assimilation; we assimilate each ensemble member with a different ${\bm{\mathit{d}}}_{j}\sim\mathcal{N}({\bm{\mathit{d}}},\bm{\mathsf{C}}_{dd})$ to avoid covariance underestimation in ensemble filters (Burgers et al., 1998); $\bm{\mathsf{K}}_{r}$ is the regularized Kalman gain matrix; and $\bm{\mathsf{J}}=\mathrm{d}\,{\bm{\mathit{b}}}/\mathrm{d}\,\bm{\mathsf{M}}{\bm{\mathit{\psi}}}$ is the Jacobian of the bias estimator. Formulae 25 generalize the analytical solutions of Nóvoa et al. (2024) to the case of biased measurement. We prescribe $\bm{\mathsf{C}}_{dd}=\bm{\mathsf{C}}_{bb}$ because the model bias is defined in the observable space. We use $\gamma$ to tune the norm of the bias in (24) (Nóvoa et al., 2024). The optimal state and parameters are
@@
$\begin{bmatrix}{\bm{\mathit{\phi}}}_{j}^{\mathrm{a}}\\[10.00002pt] {\bm{\mathit{\alpha}}}_{j}^{\mathrm{a}}\end{bmatrix}=\begin{bmatrix}{\bm{\mathit{\phi}}}_{j}^{\mathrm{f}}\\[10.00002pt] {\bm{\mathit{\alpha}}}_{j}^{\mathrm{f}}\end{bmatrix}+\overbrace{\begin{bmatrix}\bm{\mathsf{C}}_{\phi q}^{\mathrm{f}}\\[10.00002pt] \bm{\mathsf{C}}_{\alpha q}^{\mathrm{f}}\end{bmatrix}\left\{\bm{\mathsf{C}}_{dd}+(\mathbb{I}+\bm{\mathsf{J}})^{\mathrm{T}}(\mathbb{I}+\bm{\mathsf{J}})\bm{\mathsf{C}}_{qq}^{\mathrm{f}}+\gamma\bm{\mathsf{J}}^{\mathrm{T}}\bm{\mathsf{J}}\bm{\mathsf{C}}_{qq}^{\mathrm{f}}\right\}^{-1}}^{\text{Regularized Kalman gain,}\,\bm{\mathsf{K}}_{r}}\dots$
@@
$\dots\Big{[}\left(\mathbb{I}+\bm{\mathsf{J}}\right)^{\mathrm{T}}\overbrace{\left({\bm{\mathit{d}}}_{j}+{\bm{\mathit{b}}}_{d}-{\bm{\mathit{y}}}_{j}^{\mathrm{f}}\right)}^{\text{Corrected innovation}}-\gamma\bm{\mathsf{J}}^{\mathrm{T}}{\bm{\mathit{b}}}\Big{]}$ (26)
@@
The r-EnKF defines a ‘good’ analysis from a biased model if the unbiased state ${\bm{\mathit{y}}}$ matches the truth, and the model bias ${\bm{\mathit{b}}}$ is small relative to the truth. The underlying assumptions of this work are that (i) our low-order model is qualitatively accurate such that the model bias ${\bm{\mathit{b}}}$ has a small norm; and (ii) the sensors are properly calibrated. In the limiting case when the assimilation framework is unbiased (i.e., ${\bm{\mathit{b}}}={\bm{\mathit{0}}}$, and ${\bm{\mathit{b}}}_{d}={\bm{\mathit{0}}}$), the r-EnKF (3.4) becomes the bias-unregularized EnKF (A).
4 Reservoir computing for inferring model bias and measurement shift
@@
To apply the r-EnKF (3.4) we must provide an estimate of the model bias and the measurement shift. We employ an echo state network (ESN) for this task. ESNs are suitable for real-time data assimilation because (i) they are recurrent neural networks, i.e., they are designed to learn temporal dynamics in data; (ii) they are based on reservoir computing, hence they are universal approximators (Grigoryeva and Ortega, 2018); (iii) they are general nonlinear auto-regressive models (Aggarwal, 2018); and (iv) training an ESN consists of solving a linear regression problem, which provides a global minimum without back propagation. In this work, we generalize the implementation of Nóvoa et al. (2024) to account for both model biases and measurement shift (§4.1). We propose a training dataset generation based on time-series correlation (§4.2).
@@ 4.1 Echo state network architecture and state-space formulation
We propose an ESN that simultaneously estimates ${\bm{\mathit{b}}}$ and ${\bm{\mathit{b}}}_{d}$ (where ‘f’ indicates that the biases are forecast quantities, which are equivalent to the actual biases at convergence). Importantly, the biases ${\bm{\mathit{b}}}$ and ${\bm{\mathit{b}}}_{d}$ are not directly measurable in real-time, but we have access to the innovation, which is linked to the biases through (23). Once we have information on the innovation and the model bias (the measurement shift), we can estimate the measurement shift (the model bias) through Eq. 23.
@@
Figure 4a represents pictorially the proposed ESN at time $t_{k}$, with its three main components: (i) the input data, which are the mean analysis innovations, i.e, $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$; (ii) the reservoir, which is a high-dimensional state is characterized by the sparse reservoir matrix $\bm{\mathsf{W}}\in\mathbb{R}^{N_{r}\times N_{r}}$ and vector ${\bm{\mathit{r}}}_{k}\in\mathbb{R}^{N_{r}}$ ($N_{r}\gg N_{q}$ is the number of neurons in the reservoir states); and (iii) the outputs, which are the mean innovation and the model bias at the next time step, i.e., $\bar{{\bm{\mathit{i}}}}_{k+1}$ and ${\bm{\mathit{b}}}_{k+1}$. The inputs to the network are a subset of the output (i.e., the innovation only). The sparse input matrix $\bm{\mathsf{W}}_{\mathrm{in}}\in\mathbb{R}^{N_{r}\times(N_{q}+1)}$ maps the physical state into the reservoir, and the output matrix $\bm{\mathsf{W}}_{\mathrm{out}}\in\mathbb{R}^{N_{q}\times(N_{r}+1)}$ maps the reservoir state back to the physical state.
@@
Figure 4: Schematic representation of the proposed echo state network architecture for model bias and innovation prediction. (a) Compact architecture showing the open-loop and closed-loop configurations; and (b) unfolded architecture starting at an analysis step is at $t_{k}$, in which there is one open-loop step followed by a closed-loop forecast. In open loop the input to the ESN is the analysis mean innovation $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$, i.e., the difference between the raw acoustic pressure data and the analysis model estimate. The outputs from the ESN are (i) the innovation at the next time step $\bar{{\bm{\mathit{i}}}}_{k+1}$ (which becomes the input to the next time step in closed loop configuration); and (ii) the model bias ${\bm{\mathit{b}}}_{k+1}$, i.e., an estimate of the difference between the presumed acoustic state ${\bm{\mathit{d}}}_{k+1}^{\dagger}$ and the model estimate ${\bm{\mathit{q}}}_{k+1}$.
@@
Further, Figure 4a shows the two forecast settings of the echo state network: open-loop, which is performed when data on the innovations are available, and the closed-loop, in which the ESN runs autonomously using the output as the input in the next time step. Figure 4b shows the unfolded architecture starting at time $t_{k}$ when observations of the innovations are available (i.e., at the analysis step). At this time, the reservoir is re-initialized with the analysis innovations such that the input to the ESN is $\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}_{k}={\bm{\mathit{d}}}_{k}-\bm{\mathsf{M}}\bar{{\bm{\mathit{\psi}}}}_{k}^{\mathrm{a}}$. From this, the ESN estimates at the next time step $t_{k+1}$ the model bias and the innovation $\bar{{\bm{\mathit{i}}}}_{k+1}$, which is used as initial condition in the subsequent forecast step. Mathematically, the equations that forecast the model bias and mean innovation in time are
$[{\bm{\mathit{b}}}_{k+1};\bar{{\bm{\mathit{i}}}}_{k+1}]$ $=\bm{\mathsf{W}}_{\mathrm{out}}\left[{\bm{\mathit{r}}}_{k+1};1\right]$
@@
${\bm{\mathit{r}}}_{k+1}$ $=\textrm{tanh}\left(\sigma_{\mathrm{in}}\bm{\mathsf{W}}_{\mathrm{in}}\left[\bar{{\bm{\mathit{i}}}}^{\star}_{k}\odot{\bm{\mathit{g}}};\delta_{r}\right]+\rho\bm{\mathsf{W}}{\bm{\mathit{r}}}_{k}\right),$ (27)
@@
where $\bar{{\bm{\mathit{i}}}}^{\star}=\bar{{\bm{\mathit{i}}}}^{\mathrm{a}}$ in open loop and $\bar{{\bm{\mathit{i}}}}^{\star}=\bar{{\bm{\mathit{i}}}}$ in closed loop; the ${\tanh(\cdot)}$ operation is performed element-wise; the operator $\odot$ is the Hadamard product, i.e., component-wise multiplication; and ${\bm{\mathit{g}}}=[g_{1};\dots;g_{2N_{q}}]$ is the input normalizing term with $g_{q}^{-1}=\max{({\bm{\mathit{u}}}_{q})}-\min{({\bm{\mathit{u}}}_{q})}$, i.e., $g_{q}^{-1}$ is the range of the $q^{\mathrm{th}}$ component of the training data $\bm{\mathsf{U}}$; $\delta_{r}$ is a constant used for breaking the symmetry of the ESN (we set $\delta_{r}=0.1$(Huhn and Magri, 2020). We define $\bm{\mathsf{W}}_{\mathrm{in}}$ and $\bm{\mathsf{W}}$ as sparse and randomly generated, with a connectivity of 3 neurons (see Jaeger and Haas, 2004, for details). We compute the matrix $\bm{\mathsf{W}}_{\mathrm{out}}$ during the training.